# Just-in-time and distributed task representations in language models

### v2

@@ Abstract
Many of language models’ impressive capabilities originate from their in-context learning: based on instructions or examples, they can infer and perform new tasks without weight updates. In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context. We focus on “transferrable” task representations—vector representations that can restore task contexts in another instance of the model, even without the full prompt. We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context. Specifically, when more examples are provided in the context, transferrable task representations successfully condense evidence. This allows better transfer of task contexts and aligns well with the performance improvement. However, this evidence accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context. Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask. For longer and composite tasks, models rely on more temporally-distributed representations. This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process that language models use to perform new tasks on the fly.
@@ 1 Introduction
Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; Brown et al.,, 2020): the emergent ability to learn tasks from few-shot examples in context. This discovery has led to a variety of works exploring the behavioral features of ICL (e.g. Sclar et al.,, 2024; Min et al.,, 2022). Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024). The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities, leading to discoveries such as induction heads (e.g. Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g. Akyürek et al.,, 2022; Von Oswald et al.,, 2023).
@@
Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023). These task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt. This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream. Such intervention reinstates the task context and allows the model to perform the task without any explicit instructions or demonstrations. These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
@@
The discovery of transferrable task representations raises several intriguing questions about their dynamics and generality. How and when are these internal task representations formed throughout the context? How do these dynamics depend upon task complexity? A simple, intuitive hypothesis is that these representations develop gradually, e.g. at a rate that depends on task complexity. The representations might accrue evidence across tokens and refine monotonically into a more stable and robust task representation. This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
We set out to investigate these transferrable task representations to understand how the dynamics of ICL are exhibited within them. Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
@@
  • We find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
@@
  • The fleeting but transferrable task representations condense evidence from multiple examples. However, they also tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
@@
  • Models do not appear to condense task knowledge into local task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
@@
  • Finally, models can form distinct representations to support generating the same responses when solving tasks independently vs. as part of a broader context.
@@
Figure 1: Understanding how task representations develop over context. A. A schematic of extracting transferrable task representations and restoring task contexts in zero-shot settings. The highlighted tokens indicate the source and target for extracting and injecting task representations. B. Transferrable task representations restore task accuracy on zero-shot prompts. Results are aggregated over all models for simple tasks (see Appendix A). Error bars indicate the 95% CI over tasks. C. An overview of the development of different task representations over context. Solid bars: recontextualized zero-shot accuracy for task vectors extracted from different tokens. Transparent bars: task identity decoding accuracy from different token representations.
@@ 2 Methods
For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024). The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word or translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words). In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings. These include: repeating a simple task three times (e.g. antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose both the first and the last word in the list), and reversing or shifting an entire word list. Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each input item. See Appendix A for the full set of tasks.
@@
In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (Team et al.,, 2025). In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (Yang et al., 2025a, ).
@@
We primarily investigate task vectors discovered in Hendel et al., (2023) as a window to study language models’ transferrable task representations. We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query, as shown in Figure 1A. Task vectors are the layer residual activations extracted from the last token before answer generation (in the example in Figure 1A, this corresponds to the highlighted colon token). Hendel et al., (2023) showed that task vectors can reinstate task performance on a different query even without any prior context. Specifically, when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any prior few-shot examples.
We replicate and extend the procedure outlined in Hendel et al., (2023). For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting. As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations. We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers). The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task. This best layer is subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset. Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
@@
We then studied whether transferrable task representations form in tokens other than the colon token. We extracted layer residual token activations for other format tokens in the context, including the “Q”, the “:” following “Q”, the “A”, and the new-line token before “A”. We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer. For each of the non-colon tokens, we repeat the search for the layer that best captures task representations. All token representations are evaluated on the extent to which they restore task accuracy on zero-shot prompts.
Decoding task identity
We also studied whether token representations throughout the context contain robust task identity information. We trained simple linear decoders to predict the task category from the layer residual activations of different tokens. At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder. All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01. We report the decoding accuracy across the 25% held-out test representations.
@@ 3 Results
How do language models infer new tasks in-context? We leverage the transferability of task contexts to understand how models accrue evidence and refine task representations. We find that language models indeed form stronger task representations that aggregate in-context evidence. However, this evidence accrual process is non-gradual and happens in a sporadic way. In particular, models only form effective task representations that transfer task contexts at certain tokens (Figure 1C, also see Figure 2A). This is in strong contrast to a persistent sensitivity of task identity across tokens in the context (Figure 2B). We also find that the surprisingly local transferrable task representations tend to only capture a minimal “task scope.” Below, we discuss these findings in more detail.
@@
3.1 Local task representations can accrue evidence
@@
Consistent with the behavioral gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations also reflect increased task certainty with increased context (Figure 3A). Task vectors extracted following more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples. This suggests that language models condense information from multiple examples and form better task representations, even though the task representations extracted are fairly local (i.e., from a single token in the few-shot prompts).
@@
As models appear to successfully accrue evidence in these highly local representations, we sought to understand how the task representations themselves changed over more examples (Figure 4A). For this analysis, we look at the task representations extracted from the mode best layer across different tasks. This is to control for magnitude differences of the residual activations across layers and make a fair comparison. Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
In Gemma V3 models, as we increase the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot prompts. This can signal that in-context task representations tend to denoise or converge to more stable representations as models gain evidence. The magnitude (L2-norm) of the task vectors also tends to decrease over time. However, for both the variance and magnitude, there were considerable differences between tasks and model sizes. Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2). For certain tasks, the magnitude of the task representations first increases then decreases given more examples. We also note that both the variance and magnitude of task vectors tend to increase per more examples in some Qwen3 models (Figure S8), suggesting that different models may develop different strategies for refining task representations.
@@
3.2 Task representations exhibit temporal locality
The analyses above confirm that transferrable task representations benefit from increasing in-context evidence and converge to better representations. To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
@@ 3.2 Task representations exhibit temporal locality
We tested whether extracted representations from other format tokens in the prompt can also restore the corresponding task contexts. These include “Q”, the “:” following “Q”, “A”, and the new-line token before “A”, which are all shared across examples, tasks, and contexts. As before, we patched the activations at the same layer, but onto the corresponding format token instead of the last colon token in zero-shot prompts. As shown in Figure 3B, transferrable task representations generally do not form in the residual activations across layers in these tokens. This is true across the number of examples provided in the prompt, leading to the developmental trajectory of transferrable task representations shown in Figure 1C.
@@
3.3 Task representations exhibit scope locality

### v3

@@ Abstract
Many of language models’ impressive capabilities originate from their in-context learning: based on instructions or examples, they can infer and perform new tasks without weight updates. In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context. We study two different task representations: those that are “transferrable”—vector representations that can transfer task contexts to another model instance, even without the full prompt—and simpler representations of high-level task categories. We show that transferrable task representations evolve in non-monotonic and sporadic ways, while task identity representations persist throughout the context. Specifically, transferrable task representations exhibit a two-fold locality. They successfully condense evidence when more examples are provided in the context. But this evidence accrual process exhibits strong temporal locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context. In some cases, transferrable task representations also show semantic locality, capturing a small task “scope” such as an independent subtask. Language models thus represent new tasks on the fly through both an inert, sustained sensitivity to the task and an active, just-in-time representation to support inference.
@@ 1 Introduction
Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; brown2020language): the emergent ability to learn tasks from few-shot examples in context. This discovery has led to a variety of works exploring the behavioral features of ICL (e.g. sclar2024quantifying; min2022rethinking). Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (agarwal2024many; anil2024many). The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities, leading to discoveries such as induction heads (e.g. olsson2022context), the core circuits responsible for learning from in-context examples (cho2024revisiting; bakalova2025contextualize), and how ICL implicitly refines a model of in-context evidence (e.g. akyurek2022learning; von2023transformers).
@@
Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (todd2023function; hendel2023context). These task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt. This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream. Such intervention reinstates the task context and allows the model to perform the task without any explicit instructions or demonstrations. These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (davidson2025different; huang2024multimodal; luo2024vision).
@@
The discovery of transferrable task representations raises several intriguing questions about models’ representations of new tasks. How and when are transferrable task representations formed throughout the context? How do they differ from other types of task representations that models may use? A simple, intuitive hypothesis is that models develop representations for new tasks gradually. Task representations might reflect evidence accrual across tokens in the context and refine monotonically into more stable and robust representations. This view aligns with the behavioral findings that models perform better with more examples in-context (agarwal2024many; anil2024many).
We set out to investigate how the dynamics of ICL are reflected in different types of task representations. Our findings suggest a nuanced picture:
@@
  • We find a stark contrast between two different task representations: an inert representation of task identity is more continuously presented throughout the context, but transferrable task representations only activate sporadically at key tokens.
@@
  • The fleeting but transferrable task representations can support longer generation. However, their ability to guide model behavior also decays across independent subtask contexts.
@@
  • Models rely on more distributed task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
@@
  • Finally, models form distinct representations of a task when solving it independently vs. as part of a broader context.
@@
Figure 1: Understanding how task representations develop over context. A. A schematic of extracting transferrable task representations and restoring task contexts (via patching) in zero-shot settings. The highlighted tokens indicate the source and target for extracting and injecting task representations. B. Transferrable task representations restore task accuracy on zero-shot prompts. Results are aggregated over all models for simple tasks (see Appendix A). Error bars indicate the 95% CI over tasks. C. An overview of the development of different task representations over context. Solid bars: recontextualized zero-shot accuracy for task vectors extracted from different tokens. Transparent bars: task identity decoding accuracy from different token representations.
2 Related work
Since the discovery that large language models exhibit emergent in-context learning (brown2020language), there has been substantial interest in investigating this capability and its mechanistic basis. From a behavioral perspective, many subsequent works have explored how ICL could develop from implicit meta-learning of data properties (xie2022explanation; chan2022data), and how this may relate to the broader set of language model capabilities (chen2024parallel; lampinen2024broader). Some of this work has focused on the surprising fragility of ICL to subtle prompt changes (e.g. sclar2024quantifying); conversely, others have highlighted how ICL may be overly robust, allowing “learning” common tasks even if the labels are randomized (min2022rethinking). One particularly relevant focus of behavioral work on ICL has been on the dynamics of in-context learning; for example, how adding many example shots can improve performance on difficult tasks, or even those discouraged in post-training (agarwal2024many; anil2024many).
From a mechanistic perspective, olsson2022context showed in-context learning is supported by induction heads, and other work has studied how they might develop over training (edelman2024evolution; singh2025strategy). Extensive ablations have also elucidated core circuits supporting the aggregation of demonstrations from few-shot examples (bakalova2025contextualize; cho2024revisiting). Some recent work found that models may create transferrable task representations. These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance. hendel2023context demonstrated an instance of such representation: representations at intermediate layers of the last token in few-shot prompts can be injected to mimic the effect of a few-shot prompt. Concurrent work from todd2023function identified “function vectors,” which aggregate the effects of multiple attention heads to convey task information. Subsequent work has generalized these findings, exploring how function vectors can emerge from instructions (davidson2025different) and how task vectors capture task representations across modalities (huang2024multimodal; luo2024vision). Other works have explored how these representations emerge over training (dong2025understanding; yang2025task; yin2025attention), revealed their limitations (dong2025understanding; tikhonov2025one), and extended the methods to more robustly restore task contexts in zero-shot settings (li2024implicit; saglam2025learning). Building on these studies, our work characterizes the dynamics of different task representations both within examples in individual tokens and across the context, as well as across different task types. We compare and contrast these dynamics and additionally analyze how different task representations co-exist.
@@ 3 Methods
For our analyses, we built upon tasks from prior work on transferrable task representations (hendel2023context; todd2023function). The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word or translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words). In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings. These include: repeating a simple task three times (e.g. antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose both the first and the last word in the list), and reversing or shifting an entire word list. Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each input item. See Appendix A for the full set of tasks.
@@
In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (team2025gemma). In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (qwen3).
Decoding task identity
We study a simple representation for new tasks given in-context evidence: whether token representations throughout the context contain robust task identity information. We trained simple linear decoders to predict the task category from the layer residual activations of different tokens. At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder. All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01. We report the decoding accuracy across the 25% held-out test representations.
@@
We primarily investigate task vectors discovered in hendel2023context as a window to study language models’ transferrable task representations. In Appendix D, we show that alternative extraction methods of transferrable task representations like function vectors (see todd2023function) show consistent results. We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query, as shown in Figure 1A. Task vectors are the layer residual activations extracted from the last token before answer generation (in the example in Figure 1A, this corresponds to the highlighted colon token). hendel2023context showed that task vectors can reinstate task performance on a different query even without any prior context. Specifically, when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any prior few-shot examples.
We replicate and extend the procedure outlined in hendel2023context. For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting. As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations. We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; team2025gemma), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers). The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task. This best layer is subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset. Consistent with prior results, we generally find that task vectors extracted and patched at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
@@
We then repeat this intervention on other tokens to understand when transferrable task representations form. We extracted layer residual token activations for other format tokens in the context, including the “Q”, the “:” following “Q”, the “A”, and the new-line token before “A”. We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer. For each of the non-colon tokens, we repeat the search for the layer that best captures task representations. All token representations are evaluated on the extent to which they restore task accuracy on zero-shot prompts.
Lastly, we also explore a set of cross-token analyses to investigate the extent to which different tokens share representational or functional roles (see main text).
Characterizing the state of task representations.
We use specific terminology to describe the representational dynamics of the model’s internal state, distinct from the methodological operations (e.g., activation patching) used to probe them. By recontextualization, we refer to the phenomena that transferrable task representations reinstate task contexts and enable inference in a zero-shot setting. When discussing temporal properties, we refer to the distribution of representations along the sequence dimension. Finally, we use task scope to capture the semantic extent of the inference behavior enabled by recontextualization.
@@ 4 Results
How do language models represent new tasks in-context? We use the transferability of task contexts to understand when models form transferrable task representations, as opposed to maintaining a simple representation of task identity. We first show that transferrable task representations like task vectors indeed aggregate in-context evidence. We then show that this evidence accrual process happens in a sporadic way, with transferrable task representations only forming at certain tokens (Figure 1C, also see Figure 2A). This is in strong contrast to the persistence of representation for task identity across tokens in the context (Figure 2B). We also find that, in different settings, transferrable task representations can support generation for longer tasks or only support a minimal “task scope.” Below, we discuss these findings in more detail.
@@
4.1 Transferrable task representations accrue evidence
@@
Consistent with the behavioral gain from including more examples in-context (e.g., anil2024many), we confirm that transferrable task representations also reflect increased task certainty with increased context (Figure 3A). Task vectors extracted following more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples. This suggests that language models condense information from multiple examples and form better task representations, even though the task representations extracted are fairly local (i.e., from a single token in the few-shot prompts).
@@
As models appear to successfully accrue evidence in these highly local representations, we sought to understand how the task representations themselves changed over more examples (Figure 4A). We look at the task vectors extracted from the mode best layer across different tasks. This is to control for magnitude differences of the residual activations across layers and make a fair comparison. Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (hendel2023context).
In Gemma V3 models, as we increase the number of examples, we generally found reduced variance among task vectors extracted from different k-shot prompts. This can signal that in-context task representations tend to denoise or converge to more stable representations as models gain evidence. The magnitude (L2-norm) of the task vectors also tends to decrease over time. However, for both the variance and magnitude, there were considerable differences between tasks and models. Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2). For certain tasks, the magnitude of the task representations first increases then decreases given more examples. We also note that both the variance and magnitude of task vectors tend to increase per more examples in some Qwen3 models (Figure S8), suggesting that different models may develop different strategies for refining task representations.
@@
4.2 Different task representations exhibit distinct temporal profiles
The analyses above confirm that transferrable task representations benefit from increasing in-context evidence and converge to better representations. To understand the full temporal profile of this accrual process, we repeated the task vector intervention on other tokens in the prompt, which revealed that these representations do not strengthen monotonically. We show that this temporal locality is unique to transferrable task representations, as representations for task identity are widespread.
@@ 4.2 Different task representations exhibit distinct temporal profiles
We tested whether extracted representations from other format tokens can also restore the corresponding task contexts. These include “Q”, the “:” following “Q”, “A”, and the new-line token before “A”, which are all shared across examples, tasks, and contexts. As before, we patched the activations at the same layer, but onto the corresponding format token instead of the last colon token in zero-shot prompts. As shown in Figure 3B, transferrable task representations generally do not form in the residual activations across layers in these tokens. This is true across the number of examples provided in the prompt, leading to the developmental trajectory of transferrable task representations shown in Figure 1C.
@@
Non-trivial cross-token representational and functional transfer.
We primarily observe task identity representations in all tokens and transferrable task representations in the key pre-answer colon tokens. However, we note that these two types of representations are not entirely binary. We performed a set of cross-token task vector transfer experiments: both using representations from non-key tokens in few-shot prompts to intervene on the key colon token in zero-shot prompts, and using representations from the key colon token to intervene on non-key tokens. More details are presented in Appendix E. In both cases, we observe some non-trivial success of task recontextualization. For example, representations extracted from the colon following the “Q” token (prior to query presentation) can partially restore task contexts. In Qwen3 models, we also find that non-key token sites can functionally participate in task recontextualization when injected with effective transferrable task representations. These results suggest that task identity signals and transferrable task knowledge may coexist in token representations. Furthermore, token sites that do not usually form transferrable task representations may nonetheless preserve a functional role to establish task contexts.
We then further explore the overlap between identifiable representations across tokens. We show that task identity regressions fit on non-transferrable representations from non-key tokens generalize to transferrable ones, that there is partial overlap between the identifiable-non-transferrable subspaces and the transferrable ones, and that the degree of generalization and overlap is modulated by the same task features as in our other experiments. These results reinforce our claim that these different task representations have distinct dynamics and can coexist, but show additional nuance in their partial overlap.
4.3 Transferrable task representations capture variant scope locality